# Stability of Non-Hyperbolic Sets

### v1

@@ 3 The Hénon family
Consider the map $f(x)=2-x^{2}$. Then x=2x=-2 is an expanding fixed point and $f^{2}(0)=-2$ where 0 is the critical point. For given $b>0$ and $a$ large enough, the Hénon map $F_{a,b}$ has an horse-shoe. By decreasing $a$ to $a(b)$, one arrives at the first homoclinic tangency. Hence, there exists an analytic curve $b\mapsto a(b)$ with $a(0)=2$ such that the parameter $(a(b),b)$ corresponds to a Hénon map with an homoclinic tangency of the saddle point $p(b)$ which is a continuation of p(0)=(2,2)p(0)=(-2,-2). For all $b$ positive and small enough $F_{a(b),b}$ has a strong homoclinic tangency and the Hénon family $F_{a,b}$ restricted to a small neighborhood of $(a(b),b)$ is an unfolding. ∎
@@ 4 Existence of secondary tangencies
In this section we prove the existence of secondary tangencies in a certain strip of parameters that we introduce below.
Remember the choice of $\theta$ in (2.16) and (2.17). Denote by $C$ the entry of the matrix in Lemma 5 corresponding to the critical point $c$, see Definition 3. Let $\alpha\in(0,1)$ and let $n_{0}=[n\alpha]+O(1)$. Choose $\epsilon>0$, $T$ a positive large constant and define, for large $n$,
$\mathcal{B}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.\left(\frac{\epsilon}{\mu^{\frac{n_{0}}{2}}}-C\right)\lambda^{\theta n}\leq(a-a_{n}(t))\leq\frac{T}{\mu^{n+\frac{n_{0}}{2}}}\right\}.$ (4.1)
The strip $\mathcal{B}_{n}$ is built to contain the curves of tangencies. The exact choice of $n_{0}$ is made in (4.16).
Denote by $E=C-\frac{\epsilon}{\mu^{{n_{0}}/{2}}}$. Observe that in this definition of the strip, $\mu$ and $\lambda$ are taken in the parameter $(t,a_{n}(t))$. The following remark points out the distortion of the eigenvalues when changing the parameters in $\mathcal{B}_{n}$.
Remark 9.
[…]
$\left(\frac{\mu(t,a_{n}(t))}{\mu(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(n\lambda(t,a_{n}(t))^{\theta n}\right),\left(\frac{\lambda(t,a_{n}(t))}{\lambda(\tilde{t},\tilde{a})}\right)^{n}=1+O\left(n\lambda(t,a_{n}(t))^{\theta n}\right)$
where we used the fact that $\partial\mu/\partial t$ and $\partial\mu/\partial a$ are bounded. Moreover if $(t,a)\in\mathcal{B}_{n}$, then
$\frac{\mu(t,a_{n}(t))^{n}(a-a_{n}(t))}{\lambda(t,a)^{\theta n}\mu(t,a)^{n}}\geq{E},$
@@
${\mu(t,a)^{n+\frac{n_{0}}{2}}(a-a_{n}(t))}\leq{T}.$
@@ 4.1 The combinatorics
Fix $(t,a)\in\mathcal{B}_{n}$ and in the sequel when there is no possibility of confusion, we will suppress this choice in the notation, for example, $z=z_{t,a}$, $c=c_{t,a}$ (see Definition 3), $\lambda=\lambda(t,a)$ and $\mu=\mu(t,a)$.
@@
and a local unstable manifold $W^{u}_{\text{loc}}(c^{\prime})$ of diameter $0<L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$ where is a large constant that will be chosen later.
@@
Figure 3: $N+\theta n$ iterates
Lemma 6.
For $n$ large enough and for all $(t,a)\in\mathcal{B}_{n}$,
@@
$C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)\leq|z_{y}^{(1)}(t,a)-z_{y}(t,a)|\leq\left(C+{QL}\right)\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right),$
and the tangent space $T_{z^{(1)}}W^{u}_{\text{loc}}(z^{(1)})\subset\mathbb{R}^{m-1}\times\left\{0\right\}$ is horizontal. Moreover, the limit $\lim_{n\to\infty}{|z_{y}^{(1)}(t,a)-z_{y}(t,a)|}/{\lambda^{\theta n}}$ exists.
@@
$|z_{y}^{(1)}(t,a_{n}(t))-z_{y}(t,a_{n}(t))|=C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right).$ (4.2)
Proof.
Use coordinates centered at the critical point $c$ of the parameter $(t,a)$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta\tilde{x},\Delta\tilde{y})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $z$. By Lemma 5 we have
$\Delta\tilde{y}$ $=$ $C\Delta x+Q(\Delta y)^{2}+O\left(|\Delta x|^{2}+|\Delta x||\Delta y|+|\Delta y|^{3}\right).$
Because the first coordinate of $q_{1}$ is equal to $1$, see $(F4)$ and the following normalization, we have that $\Delta x=\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)$ and since $0\leq|\Delta y|\leq L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$,
$C\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right)\leq\Delta\tilde{y} \displaystyle\leq$ $\left(C+{QL}\right)\lambda^{\theta n}+O\left(\lambda^{2\theta n}\right).$
The first claim of the lemma is then proved. The fact that, in the parameter $(t,a_{n}(t))$, $\Delta y$ coincides with $c^{\prime}$, ensures that 4.2 is also true.
For the second claim of the lemma, use the coordinates centered in $c^{\prime}$. Let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta\tilde{x},\Delta\tilde{y})=F_{t,a}^{N}\left((\Delta x,\Delta y)\right)$ written in coordinates centered in $F^{N}(c^{\prime})$. Consider $W^{u}_{\text{loc}}(c^{\prime})$ as a graph on the $y$-axis and take a tangent vector $(\Delta u,\Delta v)\in T_{c^{\prime}}W^{u}_{\text{loc}}(c^{\prime})$. Define the slope-vector at $c^{\prime}$ as $s=\Delta u/\Delta v$. Then
$s=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.3)
This estimate is obtained as follows. Let $(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y})\in W^{u}_{\text{loc}}(z)$ such that $F^{\theta n}\left(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y}\right)=c^{\prime}$ and $(\Delta u_{0},\Delta v_{0})$ be a tangent vector to $W^{u}_{\text{loc}}(z)$ at the point $\left(z_{x}+\Delta z_{x},z_{y}+\Delta z_{y}\right)$. Then there exist uniform constants $K_{1},K_{2}$ such that
${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta z_{y}\leq K_{1}{\mu^{-\theta n}},$
and because of the quadratic behavior of $W^{u}_{\text{loc}}(z)$,
$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}\leq|\Delta z_{x}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}.$
In particular
$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}|\Delta u_{0}|\leq|\Delta v_{0}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}|\Delta u_{0}|.$
Observe that
@@
$s=O\left(\frac{\lambda^{\theta n}}{\mu^{\theta n}}\frac{\Delta u_{0}}{\Delta v_{0}}\right).$
The estimate for the slope follows. Use now coordinates $\left(\Delta x^{\prime},\Delta y^{\prime}\right)$ centered around $z$ with $F^{\theta n}\left(\Delta x^{\prime},\Delta y^{\prime}\right)=\left(\Delta x,\Delta y\right)$. Then there is a non zero smooth function $R$ such that
$|\Delta x^{\prime}|=R\left(\Delta y^{\prime}\right)\left(\Delta y^{\prime}\right)^{\frac{1}{2}},$
and
$\Delta y^{\prime}=O\left(\frac{1}{\mu^{\theta n}}\right).$
In particular,
$\frac{d^{2}|\Delta x^{\prime}|}{d\left(\Delta y^{\prime}\right)^{2}}=O\left(\left(\Delta y^{\prime}\right)^{-\frac{3}{2}}\right)=O\left(\mu^{\frac{3}{2}\theta n}\right).$
This implies that
$\frac{d^{2}|\Delta x|}{d\left(\Delta y\right)^{2}}=O\left(\mu^{\frac{3}{2}\theta n}\frac{\lambda^{\theta n}}{\mu^{2\theta n}}\right)=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
Hence,
$\Delta x=s\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y)^{2}.$ (4.4)
From (4.4) and (4.3) we have that $|\Delta x|=O\left(s\Delta y\right)=o\left(\Delta y\right)$. From Lemma 5,
$\Delta\tilde{x}$ $=$ $B\Delta y+O_{x}\left(\Delta y\right)(\Delta y)^{2}$
$\Delta\tilde{y}$ $=$ $s\Delta y+Q(\Delta y)^{2}+O_{y}\left(\Delta y\right)(\Delta y)^{3},$
where $B,Q\neq 0$, $O_{x}$ and $O_{y}$ are bounded differentiable functions in $\Delta y$. After differentiating one obtains
$\frac{d\tilde{y}}{d\tilde{x}}=\frac{2Q}{B}\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.5)
where we used (4.3). As consequence $z^{(1)}$ which is the point where ${d\tilde{y}}/{d\tilde{x}}=0$ corresponds to
$\Delta y=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}.$
Indeed $W^{u}_{\text{loc}}(z^{(1)})$ has a horizontal tangency. ∎

### v3

@@ 3 The Hénon family
Consider the map $f(x)=2-x^{2}$. Then x=2x=-2 is an expanding fixed point and $f^{2}(0)=-2$ where 0 is the critical point. For given $b>0$ and $a$ large enough, the Hénon map $F_{a,b}$ has an horse-shoe. By decreasing $a$ to $a(b)$, one arrives at the first homoclinic tangency. Hence, there exists an analytic curve $b\mapsto a(b)$ with $a(0)=2$ such that the parameter $(a(b),b)$ corresponds to a Hénon map with an homoclinic tangency of the saddle point $p(b)$ which is a continuation of p(0)=(2,2)p(0)=(-2,-2). For all $b$ positive and small enough $F_{a(b),b}$ has a strong homoclinic tangency and the Hénon family $F_{a,b}$ restricted to a small neighborhood of $(a(b),b)$ is an unfoldings. ∎
@@ 4 Existence of secondary tangencies
Given an unfolding we create in this section curves corresponding to maps with a so-called secondary tangency. These curves are graphs over the $t$-axes accumulating on the primary tangency curve at $a=0$. Moreover they are fully contained in a well defined strip in parameter space which is defined below. The restriction of the initial unfolding to a neighborhood of any of these curves is again an unfolding of the new secondary tangency.
Remember the choice of $\theta$ in (2.3.1) and (2.3.2). Denote by $C$ the entry of the matrix in Lemma 2.2.12 corresponding to the critical point $c$, see Definition 2.1.13. Observe that $C$, $c$, $\lambda$ and $\mu$ all depend on the parameters $(t,a)$. For simplicity of notation we omit this dependence. Let
$n_{0}^{*}=\left[\frac{\log\left(\lambda^{2\theta}\mu^{3}\right)}{2\log\mu}n\right].$ (4.0.1)
Choose $\epsilon>0$, $T$ a positive large constant and define, for large $n$,
$\mathcal{B}_{n}=\left\{(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]\left|\right.\left(\frac{\epsilon}{\mu^{\frac{n_{0}^{*}}{2}}}-C\right)\lambda^{\theta n}\leq(a-a_{n}(t))\leq\frac{T}{\mu^{n+\frac{n_{0}^{*}}{2}}}\right\}.$ (4.0.2)
The strip $\mathcal{B}_{n}$ is built to contain the curves of secondary tangencies. The exact choice of $n_{0}$ is made in (4.1.46). Denote by $E=C-\frac{\epsilon}{\mu^{{n_{0}^{*}}/{2}}}$.
Remark 4.0.3.
Observe that, if $(t,a)\in\mathcal{B}_{n}$, then
$\left(\frac{\mu(t,a)}{\mu(t,a_{n}(t))}\right)^{n}=1+O\left(n(a-a_{n}(t))\right),$
@@
$\left(\frac{\lambda(t,a)}{\lambda(t,a_{n}(t))}\right)^{n}=1+O\left(n(a-a_{n}(t))\right).$
@@ 4.1 The combinatorics
Fix $(t,a)\in\mathcal{B}_{n}$ and in the sequel when there is no possibility of confusion, we will suppress this choice in the notation, for example, $z=z_{t,a}$, $c=c_{t,a}$ (see Definition 2.1.13), $\lambda=\lambda(t,a)$ and $\mu=\mu(t,a)$.
@@
and a local unstable manifold $W^{u}_{\text{loc}}(c^{\prime})$ of vertical size $L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$ where $L$ is a large constant that will be chosen later. The curve $W^{u}_{\text{\rm loc}}(c^{\prime})$, is a graph over the $y$-axes. In the following lemma we give an expression for the curve $W^{u}_{\text{\rm loc}}(c^{\prime})$.
Lemma 4.1.1.
In a coordinate system centered in $c^{\prime}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(c^{\prime})$ is given by
$\Delta x=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\Delta y+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y)^{2},$
where $0\leq\Delta y<L^{\frac{1}{2}}\lambda^{\frac{\theta n}{2}}$, and $\Delta x=O\left(\left({\lambda^{3}}/{\mu}\right)^{\frac{\theta n}{2}}\right)$.
Proof.
Use coordinates centered at the critical point $c^{\prime}$ of the parameter $(t,a)$ and let $(\Delta x_{c^{\prime}},\Delta y_{c^{\prime}})\in W^{u}_{\text{loc}}(c^{\prime})$. Observe that $\Delta x_{c^{\prime}}$ is a function of $\Delta y_{c^{\prime}}$. First we are going to prove that
$\Delta x_{c^{\prime}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.1.2)
$\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right),$ (4.1.3)
and
$\frac{d^{2}\Delta x_{c^{\prime}}}{d\left(\Delta y_{c^{\prime}}\right)^{2}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.1.4)
Because the first coordinate of $z$ is equal to $1$, see Remark 2.1.15 , we have that
$c^{\prime}_{y}=1+O\left(\lambda^{\theta n}\right)$ (4.1.5)
and since $(t,a)\in\mathcal{B}_{n}$, then
$z_{y}\leq a_{n}(t)+\frac{T}{\mu^{n+n_{0}^{*}/2}}\leq\frac{1}{\mu^{n}}\left(1+O\left(\lambda^{n}\right)\right)+\frac{T}{\mu^{n+n_{0}^{*}/2}},$
where we used (2.2.10). Let $m$ be the highest point of $W^{u}_{\text{loc}}(c^{\prime})$. Then
$m_{y}=1+\left(L\lambda^{\theta n}\right)^{1/2}+O\left(\lambda^{\theta n}\right)=1+O\left(\lambda^{\theta n/2}\right).$
Let $Y$ be the vertical distance from $F^{-\theta n}(m)$ and $z$ and $X$ the horizontal one. Then, since $(t,a)\in\mathcal{B}_{n}$,
$Y$ $=$ $\frac{1}{\mu^{\theta n}}\left(1+O\left(\lambda^{\theta n/2}\right)\right)-z_{y}$ (4.1.6)
$\leq$ $\frac{1}{\mu^{\theta n}}\left(1+O\left(\lambda^{\theta n/2}\right)\right)-\left[\frac{1}{\mu^{n}}\left(1+O\left(\lambda^{\theta n}\right)\right)-E\lambda^{\theta n}\right]=O\left(\frac{1}{\mu^{\theta n}}\right),$
and since $W^{u}_{\text{loc}}(z)$ is essentially a parabola, $X\leq O\left({1}/{\mu^{\theta n/2}}\right)$. From this and from the fact that $z_{x}=1$, see Remark 2.1.15, formula (4.1.2) follows. Observe that, the estimate on $X$ also gives that
$c^{\prime}_{x}=\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.1.7)
To prove (4.1.3), use coordinates $\left(\Delta x_{z},\Delta y_{z}\right)\in W^{u}_{\text{\rm loc}}(z)$ centered around $z$ and coordinates $\left(\Delta x_{F^{\theta n}(z)},\Delta y_{F^{\theta n}(z)}\right)\in W^{u}_{\text{\rm loc}}(c^{\prime})$ centered around $F^{\theta n}(z)$, i.e. $F^{\theta n}\left(\Delta x_{z},\Delta y_{z}\right)=\left(\Delta x_{F^{\theta n}(z)},\Delta y_{F^{\theta n}(z)}\right)$. Let $(u,v)$ be a tangent vector to $W^{u}_{\text{loc}}(z)$ at the point $\left(\Delta x_{z},\Delta y_{z}\right)$. Then, using the same estimates as in (4.1.6), there exist uniform constants $K_{1},K_{2}$ such that
${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta y_{z}\leq K_{1}{\mu^{-\theta n}},$ (4.1.8)
and because of the quadratic behavior of $W^{u}_{\text{loc}}(z)$,
$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}\leq|\Delta x_{z}|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}.$
In particular
$K_{2}^{-1}{\mu^{-\frac{\theta n}{2}}}|u|\leq|v|\leq K_{2}{\mu^{-\frac{\theta n}{2}}}|u|.$
Observe that
$\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}=O\left(\frac{\lambda^{\theta n}}{\mu^{\theta n}}\frac{u}{v}\right).$
The estimate for the slope follows. This estimate also give a better bound for $\Delta x$, namely $\Delta x=O\left(\left({\lambda^{3}}/{\mu}\right)^{\frac{\theta n}{2}}\right)$.
We are left to prove (4.1.4). Since $W^{u}_{\text{\rm loc}}(z)$ is the image of a straight line segment, there is a non zero uniformly bounded ${\mathcal{C}}^{2}$} function $R$ such that
$\Delta x_{z}=R\left(\Delta y_{z}\right)\left(\Delta y_{z}\right)^{\frac{1}{2}},$
and by (4.1.8), there exists a uniform constant $K_{1}$, such that
${K_{1}^{-1}}{\mu^{-\theta n}}\leq\Delta y_{z}.$
In particular,
$\frac{d^{2}\Delta x_{z}}{d\left(\Delta y_{z}\right)^{2}}=O\left(\left(\Delta y_{z}\right)^{-\frac{3}{2}}\right)=O\left(\mu^{\frac{3}{2}\theta n}\right).$
By applying the linear map $F^{\theta n}$ we have
$\frac{d^{2}\Delta x_{c^{\prime}}}{d\left(\Delta y_{c^{\prime}}\right)^{2}}=\frac{d^{2}\Delta x_{F^{\theta n}(z)}}{d\left(\Delta y_{F^{\theta n}(z)}\right)^{2}}=O\left(\mu^{\frac{3}{2}\theta n}\left(\frac{\lambda}{\mu^{2}}\right)^{\theta n}\right)=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
Finally, using (4.1.3) and (4.1.4) the curve $W^{u}_{\text{\rm loc}}(c^{\prime})$ in coordinates centered around the point $c^{\prime}$ is essentially a straight line given by
$\Delta x_{c^{\prime}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\Delta y_{c^{\prime}}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)(\Delta y_{c^{\prime}})^{2}.$ (4.1.9)
@@
Figure 3: $N+\theta n+N$ iterates
Lemma 4.1.10.
For all $(t,a)\in\mathcal{B}_{n}$,
$z_{y}^{(1)}(t,a)-z_{y}(t,a)=C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
In particular, the limit $\lim_{n\to\infty}{|z_{y}^{(1)}(t,a)-z_{y}(t,a)|}/{\lambda^{\theta n}}=C$. Moreover if $(x,y)\in W^{u}_{\text{loc}}(z^{(1)})$,
$C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq y-z_{y}(t,a)\leq(C+QL)\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
Proof.
We need to determine the the point $c^{(1)}$ in $W^{u}_{\text{loc}}(c^{\prime})$ which is mapped to $z^{(1)}$. Use coordinates centered at $c^{\prime}$ of the parameter $(t,a)$ and denote by $v^{\prime}=F^{N}(c^{\prime})$. Let $(\Delta x_{c^{\prime}},\Delta y_{c^{\prime}})\in W^{u}_{\text{loc}}(c^{\prime})$ with $(\Delta x_{v^{\prime}},\Delta y_{v^{\prime}})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $v^{\prime}$. By Lemma 2.2.12 we have
$\Delta y_{v^{\prime}}$ $=$ $C_{c^{\prime}}\Delta x_{c^{\prime}}+Q_{c^{\prime}}(\Delta y_{c^{\prime}})^{2}+O\left(\Delta x_{c^{\prime}}^{2}+\Delta x_{c^{\prime}}\Delta y_{c^{\prime}}+\Delta y_{c^{\prime}}^{3}\right).$ (4.1.11)
The point $c^{(1)}$ is given by
$0=\frac{d\Delta y_{v^{\prime}}}{d\Delta y_{c^{\prime}}}.$
By taking the derivate of (4.1.11), we obtain
$0=\frac{d\Delta y_{v^{\prime}}}{d\Delta y_{c^{\prime}}}$ $=$ $C_{c^{\prime}}\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}+2Q_{c^{\prime}}\Delta y_{c^{\prime}}+O\left(2\Delta x_{c^{\prime}}\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}+\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}\Delta y_{c^{\prime}}+\Delta x_{c^{\prime}}+3\Delta y_{c^{\prime}}^{2}\right)$
$=$ $C_{c^{\prime}}\frac{d\Delta x_{c^{\prime}}}{d\Delta y_{c^{\prime}}}\left(1+o(1)\right)+2Q_{c^{\prime}}\Delta y_{c^{\prime}}\left(1+o(1)\right)+O\left(\Delta x_{c^{\prime}}\right),$
where we used Lemma 4.1.1. Using (4.1.2), (4.1.3) we get
$\Delta y_{c^{\prime}}=O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$ (4.1.12)
Use now the point $c^{(1)}$ with coordinates centered at $c$. Let $c^{(1)}=(\Delta x_{c},\Delta y_{c})\in W^{u}_{\text{loc}}(c^{\prime})$ with $z^{(1)}=(\Delta x_{z},\Delta y_{z})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $z$. Observe that $\Delta x_{c}=\Delta x_{c^{\prime}}+c^{\prime}_{x}$ and $\Delta y_{c}=\Delta y_{c^{\prime}}+c^{\prime}_{y}-1$. Hence, by (4.1.2), (4.1.7) and (4.1.5),
$\Delta x_{c}=\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\text{ and }\Delta y_{c}=O\left(\lambda^{{\theta n}}\right).$ (4.1.13)
Moreover, by Lemma 2.2.12 we have
$\Delta y_{z}$ $=$ $C\Delta x_{c}+Q(\Delta y_{c})^{2}+O\left(\Delta x_{c}^{2}+\Delta x_{c}\Delta y_{c}+\Delta y_{c}^{3}\right).$ (4.1.14)
Using (4.1.13) in (4.1.14) we have
$z_{y}^{(1)}(t,a)-z_{y}(t,a)=\Delta y_{z}=C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
@@
If $(x,y)\in W^{u}_{\text{loc}}(z^{(1)})$, using again (4.1.14), (4.1.13) and the fact that $0\leq|\Delta y_{c}|\leq(L\lambda^{\theta n})^{1/2}+O\left(\lambda^{\theta n}\right)$, we get
$C\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right)\leq y-z_{y}(t,a)\leq(C+QL)\lambda^{\theta n}+O\left(\left(\frac{\lambda^{2}}{\mu}\right)^{\frac{\theta n}{2}}\right).$
Observe that, in the last two inequalities, for the estimate of the order terms we used $(F3)$. ∎
The curve $W^{u}_{\text{\rm loc}}(z^{(1)})$ is a graph over the $x$-axes. In the following lemma we study the expression for the curve $W^{u}_{\text{\rm loc}}(z^{(1)})$.
Lemma 4.1.15.
In a coordinate system centered in $z^{(1)}$ the ${\mathcal{C}}^{4}$} curve $W^{u}_{\text{\rm loc}}(z^{(1)})$ is given by
$\Delta y=\frac{Q}{B^{2}}\left(1+O\left(\lambda^{\theta n}\right)\right)\Delta x^{2}+O\left(\Delta x^{3}\right),$
where $0\leq\Delta y<L\lambda^{\theta n}$, and $\Delta x=O\left(\lambda^{\frac{\theta n}{2}}\right)$.
Proof.
Denote by $v^{\prime}=(v^{\prime}_{x},v^{\prime}_{y})=F^{N}(c^{\prime})$ and consider coordinates $(\Delta x_{c^{\prime}},\Delta y_{c^{\prime}})\in W^{u}_{\text{loc}}(c^{\prime})$ centered at $c^{\prime}$ with $(\Delta x_{v^{\prime}},\Delta y_{v^{\prime}})\in F_{t,a}^{N}\left(W^{u}_{\text{loc}}(c^{\prime})\right)$ centered in $v^{\prime}$. Consider also coordinates $(\Delta x_{c^{(1)}},\Delta y_{c^{(1)}})\in W^{u}_{\text{loc}}(c^{\prime})$ centered at $c^{(1)}=F^{-N}\left(z^{(1)}\right)$ with $(\Delta x_{z^{(1)}},\Delta y_{z^{(1)}})\in\left(W^{u}_{\text{loc}}(z^{(1)})\right)$ centered at $z^{(1)}=(z^{(1)}_{x},z^{(1)}_{y})$. Denote by $s=\left({\lambda^{2}}/{\mu}\right)^{\frac{\theta n}{2}}$ and observe that by Lemma 4.1.1,
$\Delta x_{c^{\prime}}$ $=$ $O\left(s\right)\Delta y_{c^{\prime}}\left[1+\Delta y_{c^{\prime}}\right],$ (4.1.16)
and by construction,
$\Delta x_{v^{\prime}}$ $=$ $z^{(1)}_{x}-v^{\prime}_{x}+\Delta x_{z^{(1)}}=\Delta x_{z^{(1)}}+O\left(s\right),$ (4.1.17)
$\Delta y_{v^{\prime}}$ $=$ $z^{(1)}_{y}-v^{\prime}_{y}+\Delta y_{z^{(1)}}=\Delta y_{z^{(1)}}+O\left(s^{2}\right),$ (4.1.18)
where we also used (4.1.12). Applyng in order (4.1.17), Lemma 2.2.12 and (4.1.16) we get the following series of equalities,
$\Delta x_{z^{(1)}}$ $=$ $\Delta x_{v^{\prime}}+O\left(s\right)$
$=$ $A_{c^{\prime}}\Delta x_{c^{\prime}}+B_{c^{\prime}}\Delta y_{c^{\prime}}+O\left(\Delta x_{c^{\prime}}^{2}+\Delta x_{c^{\prime}}\Delta y_{c^{\prime}}+\Delta y_{c^{\prime}}^{2}\right)+O\left(s\right)$
$=$ $B_{c^{\prime}}\Delta y_{c^{\prime}}\left[1+O\left(s\right)+O\left(\Delta y_{c^{\prime}}\right)\right]+O(s).$
@@
$\Delta y_{c^{\prime}}$ $=$ $\frac{\Delta x_{z^{(1)}}+O(s)+O\left(\Delta y_{c^{\prime}}^{2}\right)}{B_{c^{\prime}}\left(1+O(s)\right)}$ (4.1.19)
$=$ $\frac{\Delta x_{z^{(1)}}}{B_{c^{\prime}}\left(1+O(s)\right)}+O\left(\lambda^{{\theta n}}\right),$ (4.1.20)
@@
where we used that $\Delta y_{c^{\prime}}=O\left(\lambda^{\frac{\theta n}{2}}\right)$, see Lemma 4.1.1. Applying in order (4.1.18), Lemma 2.2.12, (4.1.16) and (4.1.19) we get
$\Delta y_{z^{(1)}}$ $=$ $\Delta y_{v^{\prime}}+O\left(s^{2}\right)$
$=$ $C_{c^{\prime}}\Delta x_{c^{\prime}}+Q_{c^{\prime}}\Delta y_{c^{\prime}}^{2}+O\left(\Delta x_{c^{\prime}}^{2}+\Delta x_{c^{\prime}}\Delta y_{c^{\prime}}+\Delta y_{c^{\prime}}^{3}\right)+O\left(s^{2}\right)$
$=$ $C_{c^{\prime}}O(s)\Delta y_{c^{\prime}}\left(1+O\left(\lambda^{\frac{\theta n}{2}}\right)\right)+Q_{c^{\prime}}\Delta y_{c^{\prime}}^{2}+O\left(s\Delta y_{c^{\prime}}^{2}+\Delta y_{c^{\prime}}^{3}\right)+O\left(s^{2}\right)$
$=$ $\left[O\left(s^{2}\right)+O\left(s\lambda^{{\theta n}}\right)+O\left(\lambda^{{2\theta n}}\right)\right]+\left[O\left(s\right)+O\left(\lambda^{{\theta n}}\right)\right]\Delta x_{z^{(1)}}$
$+$ $\left[\frac{Q_{c^{\prime}}\left(1+O(s)\right)}{B_{c^{\prime}}^{2}}+O\left(\lambda^{{\theta n}}\right)\right]\Delta x_{z^{(1)}}^{2}+\left[O\left(1\right)\right]\Delta x_{z^{(1)}}^{3}.$
Observe, that this polynomial expression has a special form, the constant and linear term are zero. This is because $z^{(}1)$ is the lowest point of the curve $W^{u}_{\text{loc}}(z^{(1)})$. Hence,
$\Delta y_{z^{(1)}}$ $=$ $\frac{Q_{c^{\prime}}}{B_{c^{\prime}}^{2}}\left(1+O\left(\lambda^{{\theta n}}\right)\right)\Delta x_{z^{(1)}}^{2}+O\left(\Delta x_{z^{(1)}}^{3}\right)$
$=$ $\frac{Q}{B^{2}}\left(1+O\left(\lambda^{{\theta n}}\right)\right)\Delta x_{z^{(1)}}^{2}+O\left(\Delta x_{z^{(1)}}^{3}\right),$
where we used that, by construction, $c$ and $c^{\prime}$ are at distance to each other proportional to $\lambda^{{\theta n}}$. ∎